# Efficient Motion Sickness Assessment: Recreation of On-Road Driving on a Compact Test Track

### v1

@@ Abstract
The ability to engage in other activities during the ride is considered by consumers as one of the key reasons for the adoption of automated vehicles. However, engagement in non-driving activities will provoke occupants’ motion sickness, deteriorating their overall comfort and thereby risking acceptance of automated driving. Therefore, it is critical to extend our understanding of motion sickness and unravel the modulating factors that affect it through experiments with participants. Currently, most experiments are conducted on public roads (realistic but not reproducible) or test tracks (feasible with prototype automated vehicles). This research study develops a method to design an optimal path and speed reference to efficiently replicate on-road motion sickness exposure on a small test track. The method uses model predictive control to replicate the longitudinal and lateral accelerations collected from on-road drives on a test track of 70 m by 175 m. A within-subject experiment (47 participants) was conducted comparing the occupants’ motion sickness occurrence in test-track and on-road conditions, with the conditions being cross-randomized. The results illustrate no difference and no effect of the condition on the occurrence of the average motion sickness across the participants. Meanwhile, there is an overall correspondence of individual sickness levels between on-road and test-track. This paves the path for the employment of our method for a simpler, safer and more replicable assessment of motion sickness.
@@
The current developments of automated vehicles (AVs) are expected to change how occupants use their commute time. Questionnaires have shown that users are keen on performing non-driving related tasks (NDRTs) during their travel with AVs [1, 2]. However, 60-70 of the population is expected to be affected by motion sickness (MS) with the envisaged scenarios of AVs (limited view on the horizon, backwards seating, etc. [3, 4]). To that end, extensive research is being performed to understand the occurrence of MS and develop mitigation strategies [5]. Both require thorough human experiments representative of on-road driving to capture various scenarios and triggers for MS modeling and test the effectiveness of vehicle control strategies in mitigating motion sickness. However, on-road driving experiments on public roads are potentially unsafe and less replicable, e.g. due to traffic and other road users. Large test tracks are a viable option [6, 7]; however, they are relatively expensive, and the complexity of on-road driving is missed.
@@
This study presents a method to recreate any on-road driving scenario on a compact test track for efficient motion sickness assessment. The effectiveness of the proposed method in replicating the occurrence of on-road motion sickness has been assessed via human experiments. The participants performed a NDRT during the experiment, ensuring limited exposure to the different visual environments between on-road and test-track driving. This ensured participants’ exposure to an environment similar to that of an AV and eliminated any additional sickening effects of the sensory conflicts arising through visual cues.
@@
A novel method is proposed to recreate any on-road sickness occurrence in a compact test track, simplifying the whole test procedure for motion comfort assessment. The method uses model predictive control (MPC) to replicate the longitudinal and lateral accelerations collected from on-road drives.
Extensive within-subject human participants experiments on public roads and replication on a test track, demonstrate the efficiency and validity of our proposed method. Results prove that only a small controlled area is needed to conduct motion sickness experiments by employing the proposed method.
II Related work
In AVs, the motion planner layer is responsible for identifying the optimal trajectory and behaviour in relation to surrounding traffic and infrastructure. Due to the importance of MS in AVs, significant research has been conducted to ensure that the motion planner can derive a comfortable driving experience that can secure occupants’ engagement in NDRTs. To that end, motion planning algorithms have been developed to consider MS accumulation as an objective for determining nausea-free trajectory. A motion planning strategy has been proposed [8, 9] that minimises the Illness Rating [10] for longitudinal and lateral accelerations without compromising the journey time. The above strategy has been adopted and employed a method for sorting motion planner alternatives based on multi-criteria evaluation, also considering metrics about ride comfort, aggressive driving, riding confidence, energy efficiency, and others [11]. However, these works employed offline optimization software (GPOPs-II [12]) without real-time testing. This strategy was later modified and implemented in real-time using ForcesPro [13] and also tested with human-in-the-loop experiments, mitigating MS by even 65 [14]. To model MS more accurately within the control modules, a frequency-shaping-based motion planning algorithm is proposed [15], by minimising the motion sickness dose value (MSDV) of the frequency-weighted longitudinal and lateral accelerations. A further step in this direction is presented in [16], proposing a receding-horizon formulation for optimization-based motion planning and also using deep reinforcement learning to minimize low-frequency accelerations [17]. Despite the importance of human body motion on motion comfort [18, 19], limited studies have considered it within the control modules. In this direction, a trajectory planning algorithm has been proposed [20] incorporating passive head tilting and the subjective vertical conflict model. Based on the knowledge from these studies, this work has the objective of recreating on-road sickness exposure on test tracks.
The challenge of recreating motion sickness exposure on a compact test track is similar to the challenge of realistic motion cueing in driving simulators. In driving simulators, so-called motion filters or washout filters are routinely applied to create simulator platform motions fitting within the static and dynamic range of the simulator [21]. Such motion cueing algorithm is used [22] to test the replication of sickness exposure from an on-road drive to six degrees of freedom moving base driving simulator, where they found the MS to be more severe for the on-road scenario over the driving simulator. This shows a gap in capturing motion sickness exposure, limiting the use of driving simulators as a full replacement for on-road MS experiments. [23]. Traditional motion cueing algorithms used in driving simulators are computationally efficient but often do not fully use the available workspace. A more effective solution is the use of model predictive control. MPC can use the workspace more effectively but it is computationally demanding, making it challenging but not impossible to use MPC in real-time [24]. In this study, we use offline MPC; therefore, computational efficiency is not prohibitive. Where driving simulators typically offer motion with six degrees of freedom, we control vehicle longitudinal and lateral motion.
The goal of this study is to derive an optimal vehicle motion on a compact test track mimicking motion sickness exposure in on-road driving. We collect naturalistic driving data on public roads, use MPC-based path planning to create motions fitting a compact track and demonstrate sickness reproduction with human participants.
III Motion planner algorithm
@@ III-A Vehicle model
III-B Model Predictive Control
@@ III-C Standstill sections
III-D Objective assessment of motion sickness
@@
where is the acceleration type, either translational or rotational, while stands for the weighted accelerations in the time domain. In this work, only the lateral () and longitudinal () accelerations are considered. After multiplying each of the by appropriate factors (), they are all summed and the overall MStotal is calculated as:
where is a multiplying factor for each term. The values are used according to the guidelines in the literature [10]. For motion sickness, the weighting filters for longitudinal , lateral and rotational motion are used based on the literature [18]. In this work, we initially calculated by using all the translational and rotational accelerations, but only the longitudinal and lateral accelerations were eventually used due to their impact on motion sickness [18]. .
@@ IV-A Overview
To evaluate our algorithm’s ability to replicate the on-road sickness exposure, a within-subject experiment (i.e., all participants experienced both conditions). The condition order (on-road and test track) was cross-over randomized to be balanced equally across the participants.
@@ IV-C Participants
The study participants were selected based on age, gender, and motion sickness susceptibility. The final sample consisted of 47 participants (29 males, 17 females, 1 non-binary). The age ranged between 17 and 68 (M = 27, SD = 12.17). Susceptibility was assessed using the motion sickness susceptibility questionnaire (MSSQ-Short) and by taking into special account the answer of the item regarding the experience of sickness in cars in the past 10 years [26]. Based on the answers, participants were categorized, and the ones falling within categories C to D were invited to participate, while participants falling within category B were also invited as participants. Participants with very high susceptibility (Category E) were excluded. The overview of the recruited participants is presented in Figure 2.
General motion sickness susceptibility assessed prior to the experiment yielded a mean MSSQ-Short of 12.97 and standard deviation of 5.99. The mean MSSQ-Short across the participants translates to a 55.90 percentile, which indicates that our participants have a slightly above-average motion sickness susceptibility. The standard deviation translates to 44.96 to 65.57 percentiles, indicating that our participants cover a limited range of susceptibility.
@@ IV-D Apparatus
As a test vehicle, a Volkswagen E-Golf equipped with an experimental SAE level 3 automated driving function was used. The vehicle was equipped with vehicle data collection devices (Figure 3) with 50 Hz sampling rate.
@@ IV-E Procedure
During both conditions, the participants were instructed to focus on their non-driving related task, which was video watching. To secure participants’ constant engagement with the NDRT, videos were edited by including unexpected events about every 90 s. The participants were asked to count these events to eventually assess their engagement with the NDRT (NDRT_counts). The NDRT ensures participants’ exposure to an environment similar to that of an AV, where they expect to engage in NDRTs. Meanwhile, this eliminates any additional sickening effects of the sensory conflicts arising through visual cues. The total number of NDRT_counts was 20 and corresponded to the maximum counts in the video that could be identified during a complete ride (i.e., 25 min).
Regarding the motion sickness assessment, participants were instructed to report their sickness on the 11-point MIsery SCale (MISC) [27, 28]. During the drives, every 1 minute, a beep was played within the car to motivate the participant to verbally state their MISC level [28]. Additionally, participants were urged to self-report any changes on the scale in between minutes if the MISC rating changed since the last response was requested. Responses were recorded through the CAN system by the experimenter’s input. The drive was aborted if a MISC level of 6 was reached or if the participant requested to stop the experiment. If no measurement has been recorded in a minute, the last MISC rating before that recorded was used. MISC ratings are filled at the level of the last rating per trial for the remaining measurement increments until the end of the ride.
@@ V Results: Part A - Motion Planning
Reference data: The average path was calculated from five on-road drives performed by three safety drivers. The velocity profile resulted by averaging the longitudinal and lateral accelerations as illustrated in Figure 4. The data collected from the vehicle for the on-road reference data are filtered using a zero-phase low-pass filter (Table II), which maintains the frequencies relevant to MS and removes the noise.
@@
Generated data (motion planner): The path is generated by the MPC framework, described in Section III, by tracking the longitudinal and lateral accelerations from the on-road reference data, while adhering to the constraints of the test track. The motion planner generates the trajectory with the velocity profile assigned to it.
@@ V-B1 Time-domain analysis
The event-based time domain accelerations are presented in Figure 6 for the event that corresponds to the grey section illustrated in Figure 5(a)-5(b). This includes two roundabouts, corners of different curvatures, and braking and accelerating events.
During the experiment in the test track, the experimental vehicle’s path-tracking controller extended the standstill phases of the reference data for a longer period than the generated. This led to accumulating minor delays ( 1 s) throughout the path for the recorded test track data. To account for these delays, an event-based comparison of the time domain accelerations is conducted, where the longitudinal and lateral accelerations, and the yaw rate are synchronized based on the standstills (Figures 6(a)-6(c)).
@@
According to Figure 6(a), the motion planner’s generated longitudinal acceleration deviates from the reference at two instances, i.e., at and (Subevent 1 and 2, Figure 5(b) and 6). At these moments, the vehicle is close to the test track boundaries (Figure 5(b)), while the reference lateral acceleration is relatively low (Subevent 1) or almost zero (Subevent 2) (Figure 6(b)). As a result, the MPC determines that decelerating the vehicle, thus decreasing the velocity, is the optimal solution to achieve the desired lateral acceleration while also taking a turn to avoid the test track’s boundaries. As a result, the yaw rate (Figure 6(c)) is significantly increased compared to the on-road data, where it was almost zero. This compromise occurs because the penalty by the adaptive weights on the offset from the centre of the track is increased urging the vehicle to return to the centre of the test track, while the priority on tracking the reference accelerations decreases. Additionally, the balance of the weights on the accelerations is prioritizing the lateral acceleration more than the longitudinal (Table I), hence the lateral acceleration is tracked better than the longitudinal. This slightly higher priority on the lateral acceleration is due to the small area of the test track not allowing high speeds to be reached, thus not being able to fully track the on-road data.
@@ V-B2 Frequency-domain analysis
To further assess the tracking performance, the accelerations are compared in the frequency domain. Figure 7(a), 7(b) and 7(d) present the measured data filtered only with a moving average filter. Figure 8(a), 8(b) and 8(d) illustrate the measured data filtered also with regards to motion sickness as detailed in Section III-D.
@@
Regarding the rest of the accelerations, which were not included in the MPC framework, significant differences are identified between on-road and test-track. The vertical accelerations () illustrate significant differences (maximum 62) around the low frequencies (0 - 2 Hz), and no loss of information for higher frequencies (2 - 20 Hz). This is expected due to the different road roughness of the test-track and the on-road. The low frequency differences are more provocative for the MS occurrence, and will be explored further regarding their impact. However, their amplitude is much lower compared to the longitudinal and lateral accelerations. As far as the roll () and the pitch () accelerations are concerned, they also illustrate significant differences (maximum 47 and 62, respectively). Regarding the roll, the lower amplitude compared to the on-road scenario might be related to the lower speed applied in the test-track. Despite the lower amplitude of the roll motion, it is expected to be more frequent in the test-track due to the multiple vehicle directional changes that the vehicle required to replicate the on-road driving scenario. Meanwhile, the difference in the pitch accelerations originated from the algorithms inability to fully replicate the longitudinal dynamics. Nevertheless, these states together with the vertical accelerations are not considered and controlled with the current MPC framework. Meanwhile, they are of low amplitude and are not expected to affect the overall motion sickness. For this, further analysis is conducted by MS weighted filtering the acceleration as detailed in Section III-D. The identified differences in the raw data of the accelerations, are also illustrated when the MS-weightings are applied (Figure 8). The yaw and longitudinal accelerations (Figure 8(a)-8(d)) still illustrate a significant difference, which is decreased in amplitude due to the filters.

### v2

@@ Abstract
The ability to engage in other activities during the ride is considered by consumers as one of the key reasons for the adoption of automated vehicles. However, engagement in non-driving activities will provoke occupants’ motion sickness, deteriorating their overall comfort and thereby risking acceptance of automated driving. Therefore, it is critical to extend our understanding of motion sickness and unravel the modulating factors that affect it through experiments with participants. Currently, most experiments are conducted on public roads (realistic but not reproducible) or test tracks (feasible with prototype automated vehicles). This research study develops a method to design an optimal path and speed reference to accurately replicate on-road motion sickness exposure on a small test track. The method uses model predictive control to replicate the longitudinal and lateral accelerations collected from on-road drives on a test track of 70 m by 175 m. A within-subject experiment (47 participants) was conducted comparing the occupants’ motion sickness occurrence in test-track and on-road conditions, with the conditions being cross-randomized. The results illustrate that the subjective (reported) motion sickness is well reproduced with an insignificant reduction on the track. Meanwhile, there is an overall correspondence of individual sickness levels between on-road and test-track. This paves the path for the employment of our method for a simpler, safer and more replicable assessment of motion sickness.
@@
The current developments of automated vehicles (AVs) are expected to change how occupants use their commute time. Questionnaires have shown that users are keen on performing non-driving related tasks (NDRTs) during their travel with AVs [1, 2]. However, 60-70 of the population is expected to be affected by motion sickness (MS) with the envisaged scenarios of AVs (limited view on the horizon, backwards seating, etc. [3, 4, 5]). Experiments have shown that MS is caused in particular by low frequency motions and the absence of external (out of the window) visual information. While the physiological mechanisms underlying the development of MS remain unclear, experiments and simulations have shown that MS can be well explained by the sensory conflict theory, which states that MS results from conflict between expected and perceived motion [6]. Extensive research is being performed to understand the occurrence of MS and develop mitigation strategies [7, 8]. Both require thorough human experiments representative of on-road driving to capture various scenarios and triggers for MS modeling and to test the effectiveness of vehicle control strategies in mitigating motion sickness. However, on-road driving experiments on public roads are potentially unsafe and less replicable, e.g. due to traffic and other road users. Large test tracks are a viable option [9, 10]; however, they are relatively expensive, and the complexity of on-road driving is missed.
In AVs, the motion planner layer is responsible for identifying the optimal trajectory and behaviour in relation to surrounding traffic and infrastructure. Due to the importance of MS in AVs, significant research has been conducted to ensure that the motion planner can derive a comfortable driving experience that can secure occupants’ engagement in NDRTs. To that end, motion planning algorithms have been developed to consider MS accumulation as an objective for determining a nausea-free trajectory. A motion planning strategy has been proposed [11, 12] that minimises the Illness Rating [13] for longitudinal and lateral accelerations without compromising the journey time. The above strategy has been adopted and employed as method for sorting motion planner alternatives based on multi-criteria evaluation, also considering metrics about ride comfort, aggressive driving, riding confidence, energy efficiency, and others [14]. However, these works employed offline optimization software (GPOPs-II [15]) without real-time testing. This strategy was later modified and implemented in real-time using ForcesPro [16] and also tested with human-in-the-loop experiments, mitigating MS by even 65 [17]. To model MS more accurately within the control modules, a frequency-shaping-based motion planning algorithm is proposed [18], by minimising the motion sickness dose value (MSDV) of the frequency-weighted longitudinal and lateral accelerations. A further step in this direction is presented in [19], proposing a receding-horizon formulation for optimization-based motion planning and also using deep reinforcement learning to minimize low-frequency accelerations [20]. Despite the importance of human body motion on motion comfort [21, 5, 22], limited studies have considered it within the control modules. In this direction, a trajectory planning algorithm has been proposed [23] incorporating passive head tilting and the subjective vertical conflict model. Based on the knowledge from these studies, this work has the objective of recreating on-road sickness exposure on test tracks.
The challenge of recreating motion sickness exposure on a compact test track is similar to the challenge of realistic motion cueing in driving simulators. In driving simulators, so-called motion filters or washout filters are routinely applied to create simulator platform motions fitting within the static and dynamic range of the simulator [24]. Such a motion cueing algorithm is used [25] to test the replication of sickness exposure from an on-road drive to six degrees of freedom moving base driving simulator, where they found the MS to be more severe for the on-road scenario compared to the driving simulator. This shows a gap in capturing motion sickness exposure, limiting the use of driving simulators as a full replacement for on-road MS experiments [26]. Traditional motion cueing algorithms used in driving simulators are computationally efficient but often do not fully use the available workspace. A more effective solution is the use of model predictive control. MPC can use the workspace more effectively but it is computationally demanding, making it challenging but not impossible to use MPC in real-time [27]. In this study, we use offline MPC; therefore, computational efficiency is not prohibitive. Where driving simulators typically offer motion with six degrees of freedom, we control vehicle longitudinal and lateral motion.
@@
This study presents a method to recreate any on-road driving scenario on a compact test track (i.e., small size) for efficient and reproducible human in the loop motion sickness assessment (i.e., simplified test procedure of motion sickness assessment). The effectiveness of the proposed method in replicating the occurrence of on-road motion sickness has been assessed via human experiments. The participants performed a NDRT during the experiment, ensuring limited exposure to the different visual environments between on-road and test-track driving. Their focus on the NDRT was assessed using a performance quantification method. This ensured participants’ exposure to an environment similar to that of an AV and eliminated any additional sickening effects of the sensory conflicts arising through visual cues.
@@
A novel method is proposed to recreate any on-road sickness occurrence in a compact test track, simplifying and enhancing the efficiency of human in the loop motion comfort assessment. The method uses model predictive control (MPC) to replicate the longitudinal and lateral accelerations collected from on-road drives.
Extensive within-subject human participants experiments on public roads and replication on a test track, demonstrate the effectiveness and validity of our proposed method. Results prove that only a small controlled area is needed to conduct motion sickness experiments by employing the proposed method.
II Motion planner algorithm
@@ II-A Vehicle model
II-B Model Predictive Control
@@ II-C Standstill sections
III Objective assessment of motion sickness
@@
where is the acceleration type, either translational or rotational, while stands for the weighted accelerations in the time domain. In this work, only the lateral () and longitudinal () accelerations are considered. After multiplying each of the by appropriate factors (), they are summed and the overall MSDVtotal is calculated as:
where is a multiplying factor for each term. The values are used according to the guidelines in the literature [13]. For motion sickness, the weighting filters for longitudinal , lateral and rotational motion are used based on the literature [21]. In this work, we initially calculated by using all the translational and rotational accelerations, but only the longitudinal and lateral accelerations were eventually used due to their impact on motion sickness [21].
@@ IV-A Overview
To evaluate our algorithm’s ability to replicate the on-road sickness exposure, a within-subject experiment (i.e., all participants experienced both conditions) was conducted. The condition order (on-road and test track) was cross-over randomized to be balanced equally across the participants.
@@ IV-C Participants
The study participants were recruited based on age, gender, and motion sickness susceptibility. A wide range of ages (adults) was recruited. Motion sickness susceptibility was screened to avoid highly susceptible or insusceptible participants, and gender was used to create a gender-balanced dataset as much possible. The final sample consisted of 47 participants (29 males, 17 females, 1 non-binary). All the participants were healthy and without any vestibular dysfunction. The age ranged between 17 and 68 (M = 27, SD = 12.17). Susceptibility was assessed using the motion sickness susceptibility questionnaire (MSSQ-Short) and by taking into special account the answer of the item regarding the experience of sickness in cars in the past 10 years [29]. Based on the answers, participants were categorized, and the ones falling within categories C to D were recruited to participate, while participants falling within category B were also invited as participants. Participants with very high susceptibility (Category E) were excluded. The overview of the recruited participants is presented in Figure 2.
General motion sickness susceptibility assessed prior to the experiment yielded a mean MSSQ-Short of 12.97 and standard deviation of 5.99. The mean MSSQ-Short across the participants translates to 55.90 percentile, which indicates that our participants have a slightly above-average motion sickness susceptibility. The standard deviation translates to 44.96 to 65.57 percentiles, indicating that our participants cover a limited range of susceptibility.
@@ IV-D Apparatus
As a test vehicle, a Volkswagen E-Golf equipped with an experimental SAE level 3 automated driving function was used (Figure 3). The vehicle was equipped with vehicle data collection devices with a 50-Hz sampling rate.
@@ IV-E Procedure
During both conditions, the participants were instructed to focus on a non-driving related task, which was video watching. To secure participants’ constant engagement with the NDRT, videos were edited by including unexpected events about every 90 s. The participants were asked to count these events to eventually assess their engagement with the NDRT (NDRT_counts). The NDRT ensures participants’ exposure to an environment similar to that of an AV, where they expect to engage in NDRTs. Meanwhile, this eliminates any additional sickening effects of the sensory conflicts arising through uncontrolled out of the window visual cues. The total number of NDRT_counts was 20 and corresponded to the maximum counts in the video that could be identified during a complete ride (i.e., 25 min).
Regarding the motion sickness assessment, participants were instructed to report their sickness on the 11-point Motion Illness Symptoms Classification scale (MISC) [30]. During the drives, every 1 minute, a beep was played within the car to motivate the participant to verbally state their MISC level [30]. Additionally, participants were urged to self-report any changes on the scale in between minutes if the MISC rating changed since the last response was requested. Responses were recorded through the CAN system by the experimenter’s input. The drive was aborted if a MISC level of 6 was reached or if the participant requested to stop the experiment. If no measurement has been recorded in a minute, the last MISC rating before that recorded was used. MISC ratings are filled at the level of the last rating per trial for the remaining measurement increments until the end of the ride.
@@ V Results: Part A - Motion Planning
Reference data: The average path was calculated from five on-road drives performed by three safety drivers. The velocity profile resulted by averaging the longitudinal and lateral accelerations. The data collected from the vehicle for the on-road reference data are filtered using a zero-phase low-pass filter (Table II), which maintains the frequencies relevant to MS and removes the noise.
@@
Generated data (motion planner): The path is generated by the MPC framework, described in Section II, by tracking the longitudinal and lateral accelerations from the on-road reference data, while adhering to the constraints of the test track. The motion planner generates the trajectory with the velocity profile assigned to it.
@@ V-B1 Time-domain analysis
The event-based time domain accelerations are presented in Figure 5 for the event that corresponds to the grey section illustrated in Figure 4(a)-4(b). This includes two roundabouts, corners of different curvatures, and braking and accelerating events.
During the experiment in the test track, the experimental vehicle’s path-tracking controller extended the standstill phases of the reference data for a longer period than the generated. This led to accumulating minor delays ( 1 s) throughout the path for the recorded test track data. To account for these delays, an event-based comparison of the time domain accelerations is conducted, where the longitudinal and lateral accelerations, and the yaw rate are synchronized based on the standstills (Figures 5(a)-5(c)).
@@
According to Figure 5(a), the motion planner’s generated longitudinal acceleration deviates from the reference at two instances, i.e., at and (Subevent 1 and 2, Figure 4(b) and 5). At these moments, the vehicle is close to the test track boundaries (Figure 4(b)), while the reference lateral acceleration is relatively low (Subevent 1) or almost zero (Subevent 2) (Figure 5(b)). As a result, the MPC determines that decelerating the vehicle, thus decreasing the velocity, is the optimal solution to achieve the desired lateral acceleration while also taking a turn to avoid the test track’s boundaries. As a result, the yaw rate (Figure 5(c)) is significantly increased compared to the on-road data, where it was almost zero. This compromise occurs because the penalty by the adaptive weights on the offset from the centre of the track is increased urging the vehicle to return to the centre of the test track, while the priority on tracking the reference accelerations decreases. Additionally, the balance of the weights on the accelerations is prioritizing the lateral acceleration more than the longitudinal (Table I), hence the lateral acceleration is tracked better than the longitudinal. This slightly higher priority on the lateral acceleration is due to the small area of the test track not allowing high speeds to be reached, thus not being able to fully track the on-road data.
@@ V-B2 Frequency-domain analysis
To further assess the tracking performance, the accelerations are compared in the frequency domain. Figure 6(a), 6(b) and 6(d) present the measured data filtered only with a moving average filter. Figure 7(a), 7(b) and 7(d) illustrate the measured data filtered also with regards to motion sickness as detailed in Section III.
@@
Regarding the rest of the accelerations, which were not included in the MPC framework, significant differences are identified between on-road and test-track. The vertical accelerations () illustrate significant differences (maximum 62) around the low frequencies (0 - 2 Hz), and no loss of information for higher frequencies (2 - 20 Hz). This is expected due to the different road roughness of the test-track and the on-road. The low frequency differences are more provocative for the MS occurrence, and will be explored further regarding their impact. However, their amplitude is much lower compared to the longitudinal and lateral accelerations. As far as the roll () and the pitch () accelerations are concerned, they also illustrate significant differences (maximum 47 and 62, respectively). Regarding the roll, the lower amplitude compared to the on-road scenario might be related to the lower speed applied in the test-track. Despite the lower amplitude of the roll motion, it is expected to be more frequent in the test-track due to the multiple vehicle directional changes that the vehicle required to replicate the on-road driving scenario. Meanwhile, the difference in the pitch accelerations originated from the algorithms inability to fully replicate the longitudinal dynamics. Nevertheless, these states together with the vertical accelerations are not considered and controlled with the current MPC framework. Meanwhile, they are of low amplitude and are not expected to affect the overall motion sickness. For this, further analysis is conducted by MS weighted filtering the acceleration as detailed in Section III. The identified differences in the raw data of the accelerations, are also illustrated when the MS-weightings are applied (Figure 7). The yaw and longitudinal accelerations (Figure 7(a)-7(d)) still illustrate a significant difference, which is decreased in amplitude due to the filters.